What is the formula for the nth term for the infinite sequence:
1,5,9,13,17,21,...

Answer by palanisamy(496)

1,5,9,13,17,21,...
THIS IS AN aRITHMETIC SEQUENCE.
fIRST TERM A = 1
COMMON DIFFERENCE D = 4
nTH TERM = A+(n-1)D
= 1+(n-1)4
= 1+4n-4
= 4n-3